TOF during 2009 2010 runs G Scioli TOF



















- Slides: 19
TOF during 2009 -2010 runs G. Scioli TOF meeting – CERN 9 December 2009
Hardware status
Available sectors 18 (100%) All services installed and commissioned: Before 2009 data-taking start Gas system: OK (100%) Cooling system: OK (100%)
Low Voltage and Read. Out status
LV: 99. 6% good Total Good (%) Data Readout Module 72 72 100 Local Trigger Module 72 72 100 Clock and Pulser Distribution Module 36 36 100 672 649 97 TDC Readout Module
High Voltage status 1. 6% OFF
2009 cosmic data taking
TOF: 1592 runs for 1321 hours
TOF trigger configurations For more details see A. Silenzi’s talk
TPC-dedicated trigger configuration used since the beginning of data taking: back to back ( 3 SM) – 0 OB 3 6 5 7 7 8 9 10 11 6 5 4 3 2 6 5 4 3 2 7 1 7 7 1 1 7 7 7 0 0 8 0 8 17 17 9 16 16 10 11 11 15 15 15 12 14 12 12 14 14 12 13 14 13 13 13 6 5 7 7 8 9 10 11 4 3 2 1 6 5 4 3 2 7 1 7 1 7 7 7 0 0 8 0 8 17 17 9 16 10 16 11 15 11 12 13 14 15 1
TRD-dedicated trigger configuration (for horizontal TRD SMs): Back to back ( 3 SM) + back to back ( 1 SM) 6 7 8 5 4 3 6 2 7 1 7 0 17 9 10 16 11 12 13 14 15 8 5 4 3 2 1 7 0 17 9 10 16 11 12 13 14 15
Em. Cal-dedicated trigger configuration: back to back ( 1 SM) SM 4 and SM 5 6 7 8 5 4 3 6 2 7 1 7 0 17 9 10 16 11 12 13 14 15 8 5 4 3 2 1 7 0 17 9 10 16 11 12 13 14 15
In total: 468 M events with TOF in readout and as a trigger detector Cumulated statistics Wk 32 -33 detectors warming-up Wk 34: L 3=0. 0 T Dipole=0. 0 T TOF+TPC+ITS trg: TOF(+/-3) TOF+TPC+ITS trg: SPD Mutrg+Mutrk 33 M 23 K 75 K Wk 35 -36: L 3=0. 5 T / 0. 2 T Dipole=0. 7 T TOF+TPC+ITS++ trg: TOF(+/-3) TOF+TPC+ITS trg: SPD Mutrg+Mutrk 110 M 27 K 283 K Wk 37: L 3=0. 0 T Dipole=0. 0 T TOF+TPC+ITS++ trg: TOF(+/-3) Mutrg+Mutrk 165 M 114 K Wk 40 -41 -42 -43 -44: L 3=0. 0 T Dipole=0. 0 T TOF+TPC+ITS++ trg: TOF(+/-3) 160 M Wk 45 -46 -47: L 3=0. 2 T Dipole=0. 0 T TOF+TPC+ITS++ trg: SPD Wk 38/39: Maintenance 92 M
Run coordination
August – December 2009 shifts ALICE Shift Model: 6 -day blocks 48 hours between 2 shift blocks 132 data taking days: • 132 ON-CALL shifts (24 h each) 3 experts • 396 shifts (3 shifts of 8 h each per day) 66 blocks 27 TOF shifters (Bologna: 15, Salerno: 3, Roma: 1 ITEP: 6 and Kangnung: 2)
Planning for 2010 From 17 December 2009 to 31 January 2010: shutdown From 1 February 2010 to 30 November 2010: data taking 2010 data taking (only): • 35 weeks of p-p collisions • 5 weeks of Pb-Pb collisions 280 days • 280 ON-CALL shifts (24 h each) 4 experts • 840 shifts (3 shifts of 8 h each per day) 140 blocks 31 TOF shifters Bologna: Salerno: Centro Fermi: ITEP: Kangnung: 16 people 51. 6% 4 people 12. 9% 3 people 9. 7% 4 people 12. 9%
p-p collisions on 6 December 2009 After a long and exhausting night at 7: 23 AM on 6 December 2009 ALICE saw the first collisions with all the detectors ON
TIME BEAM MODE -1 h No beam All detectors ITS, Em. CAl, (TOF, TPC, TRD, PHOS HMPID…. ) Injection ZDC -3 min 0 SAFE STATE READY TO TAKE DATA AQUISITION (DAQ) STABLE +1 min STABLE V 0, Mu. Trg +15 min STABLE TOF, TRD … TPC half HV +25 min STABLE TPC nominal HV +40 min STABLE END STABLE PHYSICS_1: ITS, Em. Cal, PHOS, ZDC PHYSICS_1: All the detectors together PHYSICS_2: TOF, TRD, FMD, HMPID, PMD, V 0, T 0 PHYSICS_3: PHYSICS_4: TPC Mu. Trg, Mu. Trk TEST_1: TPC laser
TOF was ready at the beginning of the cosmic data taking in August …and it continues to be ready!!!